# Grammar-Based Code Representation: Is It a Worthy Pursuit for LLMs?

### v1

@@ 4.1 Large Language Models for Code
Since the release of ChatGPT-3.5 cha (2022) sparked a new wave of interest in LLMs, increasing focus has been on training and utilizing LLMs for code-related tasks. These models can be broadly categorized into two types. The first category consists of general-purpose models, which perform well in various natural language tasks, while also showing strong capabilities in code-related tasks. Examples of models in this category include ChatGPT OpenAI (2024), Gemini Reid et al. (2024), Claude Anthropic (2025), Qwen Team (2024), and DeepSeek Bi et al. (2024). The second category comprises models specifically trained on code data, including models such as CodeLlama Rozière et al. (2024), OpenCoder Huang et al. (2024), and DeepSeek-Coder Guo et al. (2024). Compared to general-purpose models, these specialized models can achieve comparable or superior performance on code-related tasks with fewer parameters and offer broader support for less common programming languages. However, regardless of whether they have been specifically trained for code-related tasks, these models represent programming languages in the same way as natural language—using token sequences. This hinders the model’s ability to recognize the inherent structural information of programming languages. Therefore, we leverage the grammar-based code representation to train GrammarCoder, which enhances the model’s ability to capture structural information inherent in programming languages.
@@ 4.2 Grammar-Based Code Representation
Many models attempt to incorporate grammar-based information into code representations Sun et al. (2019); Guo et al. (2022); Zhu et al. (2024); Sun et al. (2020); Xiong and Wang (2022); Rabinovich et al. (2017). These models have been validated on relatively small-scale models (fewer than 220M parameters), demonstrating that grammar-based representation helps prevent syntax errors and enhances code generation performance. For example, GrammarT5 Zhu et al. (2024) is a pre-trained model based on grammatical rules. It is trained based on CodeT5 (220M) Wang et al. (2021b) with an encoder-decoder architecture using the same training data, demonstrating that grammar-based representations can enhance model performance. However, with the emergence of LLMs, models’ size has expanded rapidly, and decoder-only architectures have gradually become mainstream. It’s unclear whether grammar-based representations remain effective in larger-scale (e.g., billion-size) decoder-only models. Moreover, beyond preventing grammatical errors, it remains unclear whether grammar-based representations provide any additional benefits. Therefore, we bridge these gaps by training and evaluating grammar-based representations in billion-scale decoder-only models. Additionally, we explore why grammar-based representation remains effective when syntax errors are rare in LLMs, providing insights into its broader impact on model performance.
@@ A.1 Mode Configuration
Config DeepSeek-Coder Qwen2.5
# parameters 1.3 B 1.5 B
# hidden_layer 24 28
# hidden_size 2,048 1,537
# intermediate_size 5,504 8,960
# attention_head 16 12
# vocabulary 32,256 151,936
Table 4: The main configuration of two different base models.
Table 4 presents the key configurations of the base models used in our study: DeepSeek-Coder and Qwen2.5. While both models are billion-scale in terms of parameter count, they exhibit differences in architectural details, particularly in vocabulary size. DeepSeek-Coder has a vocabulary size of 32,256, whereas Qwen2.5 employs a significantly larger vocabulary of 151,936. Since grammar-based representations restructure code at a syntactic level rather than relying solely on the token level, their effectiveness is not dependent on the original vocabulary. Therefore, this difference in vocabulary size can underscore the robustness of our grammar-based code representation. After incorporating grammar rules, our vocabulary sizes expand to 33,465 for DeepSeek-Coder and 153,108 for Qwen2.5. If GrammarCoder demonstrates improved performance across both base models, it would further indicate that grammar-based approaches are adaptable to different model architectures and tokenization strategies.
@@ A.3 Training Datasets and Filter
Table 5: Open-source instruction datasets are used in our instruction tuning process.
@@
For training the base models, we primarily use high-quality Python code, aligning with our focus on grammar-based code representation. Our dataset is composed of two key sources. First, we sample 10B tokens from TheStackV2 Lozhkov et al. (2024), a large-scale code corpus that provides diverse and high-quality programming samples across various domains, ensuring a strong foundation in general coding patterns and structures. Second, inspired by previous studies Huang et al. (2024); Yang et al. (2024), we incorporate 0.5B tokens of self-contained code textbooks from open-source repositories Huang et al. (2024). Unlike context-dependent snippets, these samples consist of independent tasks and corresponding code snippets, helping the model learn to generate independent and coherent programs, bridging the gap between standard pre-training and real-world interactive programming scenarios.
For training the instruct models, we use instruction data consisting of two main sources: publicly available instruction datasets and synthetically generated instruction data. Table 5 lists the open-source instruct datasets used in our training Hu et al. (2024); Huang et al. (2024); Computations (2024); Zheng et al. (2024); TokenBender (2024), each contributing to the diversity and quality of instruction tuning. All of the datasets have a permissive license for the training LLM. For synthetic instruct data, we use LLaMA3.1-70B as the base model to generate high-quality data, leveraging OSS-Instruct Wei et al. (2024b) and SelfCodeAlign Wei et al. (2024a) as synthesis methods. This approach enables us to create a large-scale instruct dataset totaling 5B tokens, further enhancing the model’s ability to follow instructions effectively.
To ensure data quality, we apply data filtering for both base models and instruct models, primarily focusing on deduplication and syntax validation. Deduplication is performed through string-based text matching to eliminate redundant samples. For syntax validation, we use Tree-sitter TreeSitter (2024) to check whether the code can be parsed into a valid syntax tree; if parsing fails, the sample is removed. These filtering steps help maintain a high-quality and diverse instruction dataset for training.
@@ B.1 Distribution of Edit Distance
Figure 3 shows the edit distance distribution for error-correct code pairs with small edit distances (less than 50, accounting for 91.18% of the test set) under different code representation approaches.
@@ B.2 Analysis of Model Outputs
While we have examined the differences between representations on existing datasets, it is also crucial to analyze whether grammar-based representation amplifies token-level subtle differences in the model’s generated outputs. Therefore, we further analyzed the inference results of Meta-Llama-3.1-70B Dubey et al. (2024) on the CodeNanoFix dataset, focusing on the edit distance between correct and incorrect code samples for the same data samples. The results show that in 25.56% of the samples, the token-level edit distance between incorrect and correct code is relatively small (less than 50). Among these samples, the average edit distance for token-based representations is 28.04, whereas for grammar-based representations, it increases to 44.56. These findings suggest that even for a 70B-scale model, generating the correct code remains challenging when token-level differences are minimal. Relying solely on token-level information may not be sufficient to distinguish critical semantic differences in code. In contrast, grammar-based representations provide additional structural information, helping the model better differentiate between similar yet semantically distinct code snippets.

### v2

@@ 4.1 Large Language Models for Code
Since the release of ChatGPT-3.5 chatgpt3.5 sparked a new wave of interest in LLMs, increasing focus has been on training and utilizing LLMs for code-related tasks chen2025sesurvey; wang2024prompt. These models can be broadly categorized into two types. The first category consists of general-purpose models, which perform well in various natural language tasks, while also showing strong capabilities in code-related tasks. Examples of models in this category include ChatGPT chatgpt2024, Gemini gemini, Claude claude, Qwen qwen2.5, and DeepSeek bi2024deepseek. The second category comprises models specifically trained on code data, including models such as CodeLlama codellama, OpenCoder opencoder, and DeepSeek-Coder guo2024deepseekcoder. Compared to general-purpose models, these specialized models can achieve comparable or superior performance on code-related tasks with fewer parameters and offer broader support for less common programming languages. However, regardless of whether they have been specifically trained for code-related tasks, these models represent programming languages in the same way as natural language—using token sequences. This hinders the model’s ability to recognize the inherent structural information of programming languages. Therefore, we leverage the grammar-based code representation to train GrammarCoder, which enhances the model’s ability to capture structural information inherent in programming languages.
@@ 4.2 Grammar-Based Code Representation
Many models attempt to incorporate grammar-based information into code representations sun2019grammar; liangbipartite; unixcoder; grammart5; treegen; l2s; abstract. These models have been validated on relatively small-scale models (fewer than 220M parameters), demonstrating that grammar-based representation helps prevent syntax errors and enhances code generation performance. For example, GrammarT5 grammart5 is a pre-trained model based on grammatical rules. It is trained based on CodeT5 (220M) codet52021 with an encoder-decoder architecture using the same training data, demonstrating that grammar-based representations can enhance model performance. However, with the emergence of LLMs, models’ size has expanded rapidly, and decoder-only architectures have gradually become mainstream. It’s unclear whether grammar-based representations remain effective in larger-scale (e.g., billion-size) decoder-only models. Moreover, beyond preventing grammatical errors, it remains unclear whether grammar-based representations provide any additional benefits. Therefore, we bridge these gaps by training and evaluating grammar-based representations in billion-scale decoder-only models. Additionally, we explore why grammar-based representation remains effective when syntax errors are rare in LLMs, providing insights into its broader impact on model performance.
@@ 5 Conclusion
6 Acknowledgment
This work is sponsored by the National Key Research and Development Program of China under Grant No. 2023YFB4503803, and the National Natural Science Foundation of China under Grant No. W2411051, No. 62232003, and No. 62402482. Kuaishou Technology also supports this work.
7 Limitations
@@ A.1 Mode Configuration
Config DeepSeek-Coder Qwen2.5 Qwen2.5
# parameters 1.3 B 1.5 B 7 B
# hidden_layer 24 28 28
# hidden_size 2,048 1,537 3,584
# intermediate_size 5,504 8,960 18,944
# attention_head 16 12 28
# vocabulary 32,256 151,936 152,064
Table 4: The main configuration of base models.
Table 4 presents the key configurations of the base models used in our study: DeepSeek-Coder, Qwen2.5-1.5B, and Qwen2.5-7B. While these models are billion-scale in terms of parameter count, they exhibit differences in architectural details, particularly in vocabulary size. DeepSeek-Coder uses a vocabulary of 32,256 tokens, whereas Qwen2.5 adopts a significantly larger vocabulary of over 150,000 tokens. Since grammar-based representations restructure code at a syntactic level rather than relying solely on the token level, their effectiveness is not dependent on the original vocabulary. Therefore, this difference in vocabulary size can underscore the robustness of our grammar-based code representation. After incorporating grammar rules, our vocabulary sizes expand to 33,465 for DeepSeek-Coder, 153,108 for Qwen2.5-1.5B, and 154680 for Qwen2.5-7B. If GrammarCoder demonstrates improved performance across both base models, it would further indicate that grammar-based approaches are adaptable to different model architectures and tokenization strategies.
@@ A.3 Training Datasets and Filter
Table 5: Open-source instruction datasets.
@@
For training the base models, we primarily use high-quality Python code, aligning with our focus on grammar-based code representation. Our dataset is composed of two key sources. First, we sample 10B tokens from TheStackV2 starcoder2, a large-scale code corpus that provides diverse and high-quality programming samples across various domains, ensuring a strong foundation in general coding patterns and structures. Second, inspired by previous studies opencoder; qwen2, we incorporate 0.5B tokens of self-contained code textbooks from open-source repositories opencoder. Unlike context-dependent snippets, these samples consist of independent tasks and corresponding code snippets, helping the model learn to generate independent and coherent programs, bridging the gap between standard pre-training and real-world interactive programming scenarios.
For training the instruct models, we use instruction data consisting of two main sources: publicly available instruction datasets and synthetically generated instruction data. Table 5 lists the open-source instruct datasets used in our training hu2024minicpm; opencoder; codesharegpt; zheng2024codefeedback; codealpaca, each contributing to the diversity and quality of instruction tuning. All of the datasets have a permissive license for the training LLM. For synthetic instruct data, we use LLaMA3.1-70B as the base model to generate high-quality data, leveraging OSS-Instruct wei2024magicoder and SelfCodeAlign wei2024selfcodealign as synthesis methods. This approach enables us to create a large-scale instruct dataset totaling 5B tokens, further enhancing the model’s ability to follow instructions effectively.
To ensure data quality, we apply data filtering for both base models and instruct models, primarily focusing on deduplication and syntax validation. Deduplication is performed through string-based text matching to eliminate redundant samples. For syntax validation, we use Tree-sitter treesitter to check whether the code can be parsed into a valid syntax tree; if parsing fails, the sample is removed. These filtering steps help maintain a high-quality dataset for training.
@@ B.1 Distribution of Edit Distance
Figure 3 shows the edit distance distribution for error-correct code pairs with small edit distances (less than 50, accounting for 91.18% of the test set) under different representation approaches.
@@ B.2 Analysis of Model Outputs
While we have examined the differences between representations on existing datasets, it is also crucial to analyze whether grammar-based representation amplifies token-level subtle differences in the model’s generated outputs. Therefore, we further analyzed the inference results of Meta-Llama-3.1-70B llama3 on the CodeNanoFix dataset, focusing on the edit distance between correct and incorrect code samples for the same data samples. The results show that in 25.56% of the samples, the token-level edit distance between incorrect and correct code is relatively small (less than 50). Among these samples, the average edit distance for token-based representations is 28.04, whereas for grammar-based representations, it increases to 44.56. These findings suggest that even for a 70B-scale model, generating the correct code remains challenging when token-level differences are minimal. Relying solely on token-level information may not be sufficient to distinguish critical semantic differences in code. In contrast, grammar-based representations provide additional structural information, helping the model better differentiate between similar yet semantically distinct code snippets.